OpenStudy (anonymous):

In a random sample of 75 individuals, 52 people said they prefer coffee to tea. 99.7% of the population mean is between _% and _%. (Choose the option closest to your answer.) the first options are 84, 80, 74 the second options are 64, 58, 54

OpenStudy (amistre64):

and what is your attempt at a confidence interval formula?

OpenStudy (amistre64):

mine is: \[\frac Xn\pm z\sqrt{\frac{1}{n}\frac{X}{n}\frac{(n-X)}{n}}\]